OpenStudy (anonymous):

0. Let f (x) = x^4 : (a) Find a linear approximation of f (x) at a = 1 (b)If dx = x = 0.25; what are dy and deltay?

OpenStudy (anonymous):

for a i got 4x-3 but not sure if its suppose to be an equation?

OpenStudy (anonymous):

and b i know dy=1 but will delta x also equal 1?

OpenStudy (phi):

f(x) = x^4 Yes, part a looks good. at x= 1, f(1) = 1^4 =1, and (1,1) is on the curve you want the equation of a line through that point, with a slope df/dx evalutated at x=1 the equation of the line using point-slope form is y - y0 = m(x- x0) with m= df/dx, the equation through (1,1) is y - 1 = m(x - 1) or y = mx -m+1 df/dx = 4x^3 . At x=1 df/dx = 4 y = 4x -4 +1 or y = 4x -3 Here is a plot:

OpenStudy (anonymous):

okay thanks..DO you know how to solve part B?

OpenStudy (phi):

I think \[ \Delta y = \frac{dy}{dx} \Delta x \]

OpenStudy (anonymous):

do you think i could solve delta y by f(1.25)-f(1) or would i have to use an equation?

OpenStudy (phi):

where did you get 1.25? (b) says x = 0.25 (and Δx= 0.25)

OpenStudy (phi):

If x= 1/4 and Δx= 1/4 I would do this dy/dx = 4 x^3 evaluated at x= 1/4--> dy/dx = 1/16 Δy = 1/16* Δx = 1/16 * 1/4 = 1/64

OpenStudy (anonymous):

okay i see read the problem wrong thanks